When i use the solve function on my ti-84 calc, and try to solve a specific equation log13(8-5x)=log13(2x) (base is 13) , it says bad guess on my calc, i tried to change the domain, but it still said bad guesss, helppp?
Answer by rothauserc(4718)

log13(8-5x)=log13(2x) (base is 13)
we can solve this for x without a calculator
from definition of logarithm, x = b^y, where y = logb(x), soooooooo
(8-5x) = 13^y and (2x) = 13^y'
but y = y', so we have
8-5x = 2x and
7x = 8 and
x = 8/7 = 1.14